A friend of yours recently graduated from law school with $125,000 in student debt. The interest rate on these loans is 6.25%. Your friend has various options on how to repay the loans and is asking for your help to understand the pros and cons of the various options.

2.    The second option is to repay the loan with a fixed amount per month for 25-years

a.    How much will your friend pay each month?

b.    How much will your friend pay in interest over the life of the loan?

 3.    The third and final option is to repay the loan using a graduated repayment plan. Under the graduated repayment plan your friend’s payments will start low and increase each month at an annualized rate of 3%. The graduated repayment plan will repay the loan over a 10-year period.

a.    How much will your friend pay the first month?

b.    How much will your friend pay the second month?

c.    How much will your friend pay the final month?

d.    How much will your friend pay in interest over the life of the loan?






1
Expert's answer
2022-03-04T08:29:53-0500

2.

a.


"A=125000(1+0.005208)=125651.4"


b."651,4\times300-125000=70720"


3.

a.

"A=125 000(\frac{0.005}{1-(1+0.005)^{-300}})=805.37"


"A=125 000(1+0.005208)=125651.4"

125651.4-125000=651.4

b.

"i=6.5+3=9.5"


"A=125 000(1+0.005+0.0025)=125963.5"


125963.5-125000=963.5


с



d.

"A=125 000(1+0.005+0.0025...)=145572.92"

145572.92-125000=20572.92